# Understanding Ethical Practices in AI: Insights from a Cross-Role, Cross-Region Survey of AI Development Teams

### v1

@@ 5.3.3. Familiarity with AI Governance Initiatives
We asked all participants to rate their familiarity with common AI ethics initiatives (see Figure 7). Between $\sim$58% and $\sim$62% of Group B participants are at least “somewhat familiar” with all initiatives. Group A participants scored ¿25% lower than their Group B counterparts. The initiatives with the largest disparities between the two groups are the “NIST Technical AI Standards” and the “NIST AI Risk Management Framework” (with a difference of 39.7% and 35.9%, respectively). Our results show that participants are generally less familiar with the initiatives than with the principles.
(a) Group A - Familiarity with AI Governance Initiatives
(b) Group B - Familiarity with AI Governance Initiatives
Figure 7. Degree of Familiarity with AI Governance Initiatives
Table 8. P-Values for Hypothesis H3a to H3e
@@
As shown in Figure 8, Group B participants in AM roles are most familiar with all initiatives, while those in QA and ISec roles are the least familiar. In addition, those who work in North America are more familiar with all initiatives than other participants, except for the EU’s AI Act, with which participants from the EU, EEA, and UK are most familiar. A similar trend appears across company types and locations: government employees are the most familiar with all AI ethics initiatives except the EU’s AI Act, for which research and academic employees demonstrate the highest familiarity. Similarly to the principles, participants from companies with 1-5 and 100+ employees are the least familiar with all initiatives. We also evaluated the correlation between familiarity and role (H3a), location (H3b), company type (H3c), company size H3d), and gender (H3e). As shown in Table 8, we did not find any correlation for the participants in Group A, with the Bonferroni correction. Without the correction, there is a correlation between the location and the average familiarity for Group A participants. However, for Group B participants, we found significant correlations between familiarity with AI governance initiatives and participants’ roles, location, company size, and gender (pvalues=0.0008,0.0003,0.000,p-values=0.0008,0.0003,0.000, and $0.0001$, respectively), both with and without the Bonferroni correction. No significant correlation was observed with company type.
Figure 8. Group B - Average Familiarity With AI Governance Initatives by Demographics
5.3.4. Perceptions of AI Governance Initiatives
@@
We evaluated whether participants perceived AI governance initiatives as having a negative impact on their work. 48%, 25%, and 15% of participants responded no, yes, and unsure, respectively. The remaining participants chose not to respond. Additionally, we examined how they perceive the impact of these initiatives and regulations in shaping the future of technology in their field (see Figure 9). The participants could respond on a Likert Scale with an option to describe the impact in more detail. The majority (66.5%) sees at least a somewhat positive impact of the regulations on the future of technology. However, many participants were uncertain or neutral (23.1%). About 10% consider a negative impact. One participant noted: “We already have that in Europe. I’m all hears for data and privacy protection laws, but sometimes these laws are maded by people outside of the industry and this has had an negative impact on the European Tech Market before.” About half of those who responded yes are mostly concerned that regulations may limit innovation. One participant remarked: “I believe that regulations for mitigating AI ethics risks might initially pose challenges, such as increased compliance costs and slower innovation. However, they can ultimately benefit the industry by building trust, ensuring fair practices, and fostering a more responsible and sustainable approach to AI development.” Others are concerned about operational challenges and costs associated with regulations. According to one participant: “If the company fails to comply with the regulations, it may face legal action, fines or other legal consequences. This not only causes direct financial losses, but may also damage the company’s reputation.” Among those who do not believe regulation negatively impacts their work, a common theme is the belief that regulation can enhance trust across the AI ecosystem. In the words of one participant: “As said, we need to ensure the data protection before using AI for our daily business. Having a higher authority to regulate that safety will increase our and our client’s trust in the systems.” Among those who are unsure, we observe a balance between recognizing the positive aspects of regulations and concerns about limiting innovation and competitiveness. One respondent commented: ”Regulations for mitigating AI ethics risks might increase compliance costs and slow innovation, potentially creating competitive disadvantages for smaller companies. However its also important to maintain certain standards.”
@@
Overall, the average familiarity with the AI governance initiatives is lower than the familiarity with the principles among various roles and demographics. However, most AI development teams have a positive outlook on how regulations will impact the future of AI development and their companies. The field is expected to benefit from increased user trust fostered by regulations and their associated safety guidelines. However, many express concern about the effects of these regulations and the motivations behind their creation, particularly regarding the challenges of implementing guidelines into AI technology which may result in slowing innovation and reducing competitiveness.
Figure 9. Impact of Future AI Regulations - All Participants
5.4. Risk Mitigation Strategies
@@
In RQ4, our goal is to understand the actual practices employed by members of AI development teams to mitigate AI ethics risks, as previously discussed. We tailored questions on AI ethics risk mitigation strategies to each role within Group B, as detailed in Table 1.
@@
Figure 10. Group B - Risk Mitigation Methods Used
5.4.1. AM Role - Mitigation Strategies
Group B participants who selected AM roles were asked how effectively they believe AI ethics principles are integrated into their products. The majority (57.38%) claimed this to be “Extremely” and “Moderately Effective,” while only 13.12% answered “Moderately Ineffective” or worse.
Table 9. Effectiveness of AI Ethics Principles Integration
Response Option Count Percentage
Extremely Effective 17 27.87%
Moderately Effective 18 29.51%
Somewhat Effective 17 27.87%
Moderately Ineffective 5 8.20%
Extremely Ineffective 3 4.92%
Prefer not to say 1 1.64%
@@ 5.4.1. AM Role - Mitigation Strategies
Table 10. P-Value and Chi-Square Statistic for Hypotheses H4a to H4e
H4a H4b H4c H4d H4e
@@
We followed up by asking those in AM roles to identify the actions their company is taking to promote AI ethics. Table 11 shows the results. “Providing AI ethics training for employees” and “implementing ethics guidelines for AI development and deployment” are considered key actions. However, only $\sim$50% of the AM roles conduct regular audits of the AI systems for ethics compliance, which is concerning given that one of the mitigation strategies in OWASP Top 10 for LLMs and AI models (OWASP GenAI Working Group, nd) is to “implement robust logging and monitoring systems to detect unusual patterns.”
Breaking down these responses by location, we observe that participants from North America are most likely to establish an AI ethics committee (57.14%), while those from the EU+EEA+UK are more likely to implement ethics guidelines (63.64%). In contrast, participants outside of these two regions are more likely to take other actions, e.g., “conduct regular audit AI systems for ethical compliance” (86.7%), “provide AI ethics training for employees” (73.3%), and/or “engage with external experts on AI ethics” (73.3%).
We proceeded by asking AMs to select who is responsible for ensuring AI ethics principles are followed. The results are shown in Table 12 (note that participants were allowed to select multiple options): 63.93% pointed to the AI development team and upper management, while only 42.62% pointed to the internal legal and compliance team or the AI ethics committee. Previous research has highlighted the critical role of dedicated teams in ensuring software/AI compliance (Prybylo et al., 2024). Other studies have emphasized the challenges of enforcing accountability for AI ethics and regulatory standards and the importance of collaborating with policymakers to implement them (Orr and Davis, 2020).
Table 11. AM Roles: Actions Taken at Company to Promote AI Ethics.
@@
Lastly, we asked participants to provide examples of how they ensure their team follows AI ethics principles in development and to describe how they promote and enforce these principles within their organization. Four of the six responses mentioned “developing ethical guidelines,” while one participant mentioned they have been “developing AI-specific red-teams.” Although the responses about promoting and enforcing the principles varied, we observed that 30.3%, 21.21%, and 21.21% of participants mentioned implementing ethical guidelines, communicating AI guidelines with employees, and educating employees on AI ethics, respectively. Of those who mentioned communicating with employees, participants highlighted, among others: “understanding and teaching staff that AI is there to aid the system, not BE the system” and “use of stringent measures on failure to adopt the principles.”
We also gained insights into some of the challenges faced during this process. For instance, one participant noted: “our role mainly revolves around building what the client requires, and we cannot control the end usage.” Another added, “we are a small company still with less than 5 workers so to keep the information up to date is easy right now.”
Table 12. AM Roles: Primary Responsibility for Ensuring AI Ethics Principles Are Followed
@@
5.4.2. RA Role - Mitigation Strategies
Group B participants in RA roles were asked how frequently AI ethics principles are included in their requirements documentation. Interestingly, fewer than 25% of participants rarely or never consider ethics in their requirements documentation (see Table 13). Similarly, about 30% believe that AI ethics requirements hinder, to some degree, the project’s outcome ( see Table 14). Regarding whether their company conducts (ethics-related) risk and impact assessments of their products, about 53% of responses were positive, with one participant noting that “another unit is responsible.” The rest responded “no” or “unsure.”

### v2

@@ 4.3.3 Familiarity with AI Governance Initiatives
We asked all participants to rate their familiarity with common AI ethics initiatives (see Figure 7). Between $\sim$58% and $\sim$62% of Group B participants are at least “somewhat familiar” with all initiatives. Group A scored $25\%$ lower than their Group B counterparts. The initiatives with the largest disparities between the two groups are the “NIST Technical AI Standards” and the “NIST AI Risk Management Framework” (with a difference of 39.7% and 35.9%, respectively). Our results show that participants are generally less familiar with the initiatives than with the principles.
(a) Group A - Degree of Familiarity
(b) Group B - Degree of Familiarity
Figure 7: Degree of Familiarity with AI Governance Initiatives
Table 9: P-Values for Hypotheses H3a to H3e
@@
As shown in Figure 8, Group B participants in AM roles are most familiar with all initiatives, while those in QA and ISec roles are the least familiar. Those who work in North America are more familiar with all initiatives than other participants, except for the EU’s AI Act, with which participants from the EU, EEA, and UK are most familiar. A similar trend appears across company types: government employees are the most familiar with all AI ethics initiatives except the EU’s AI Act, for which research and academic employees demonstrate the highest familiarity. Similarly to the principles, participants from companies with 1-5 and 100+ employees are the least familiar with all initiatives. We also evaluated the correlation between familiarity and role (H3a), location (H3b), company type (H3c), company size H3d), and gender (H3e). As shown in Table 9, we did not find any correlation for the participants in Group A, with the Bonferroni correction. Without the correction, there is a correlation between the location and the average familiarity for Group A participants. However, for Group B, we found significant correlations between familiarity with AI governance initiatives and participants’ roles, location, company size, and gender (pvalues=0.0008,0.0003,0.000,p-values=0.0008,0.0003,0.000, and $0.0001$, respectively), both with and without the Bonferroni correction. No significant correlation was observed with company type.
Figure 8: Group B - Familiarity with AI Governance Initiatives by Demographics
We conducted a LASSO analysis to identify the predictive factors within demographics, perception, and practice most strongly associated with familiarity with AI governance initiatives. For Group B participants, our model indicates that greater familiarity with AI governance initiatives is most strongly associated with the use of specific AI tools, particularly Google Cloud AI (+0.251) and LangChain (+0.240) (see Table 10). Participants who used Neural Auto Encoders (+0.192) and those who developed contingency plans for ethical risk mitigation (+0.181) also exhibited higher familiarity with AI governance initiatives. Familiarity with the Harm Prevention and Beneficence principle showed a marginally significant positive association (+0.221, p=0.054). In contrast, participants with only a high school degree demonstrated lower levels of familiarity (-0.213).
Table 10: LASSO Results: Predictors of AI Governance Initiative Familiarity
Predictor Coefficient Std. Error
AI Tools/Technologies
   Google Cloud AI $0.251^{**}$ $(0.090)$
   LangChain $0.240^{**}$ $(0.091)$
   Neural Auto Encoder $0.192^{*}$ $(0.084)$
Familiarity with AI Ethics Principles
   Harm Prevention and Beneficence $0.221$ $(0.113)$
   Accountability and Responsibility $0.178$ $(0.135)$
   Environment and Social Responsibility $0.175$ $(0.139)$
   Average principle familiarity $0.056$ $(0.212)$
Mitigation Strategies
   Create contingency plans $0.181^{*}$ $(0.090)$
   Implement transparent/explainable approaches $0.155$ $(0.087)$
Demographics
   High school degree (vs. higher education) $0.213^{*}$ $(0.085)$
Overall Mean 2.86
Observations 141
R² (test) 0.41
Features Selected 10 / 142
Note: $^{*}p<0.05$; $^{**}p<0.01$; $^{***}p<0.001$. Dependent variable: Average familiarity with 8 AI governance initiatives (1–5 scale). Coefficients are standardized (per 1-SD change in predictor).
Table 11: Impact of Future AI Regulations - All Participants
Highly +ive Somewhat +ive Neutral Somewhat -ive Highly -ive
30.90% 35.60% 23.10% 6.40% 4.00%
4.3.4 Perceptions of AI Governance Initiatives
@@
We evaluated whether participants perceived AI governance initiatives as having a negative impact on their work. 48%, 25%, and 15% of participants responded no, yes, and unsure, respectively. The remaining participants chose not to respond. We, then,examined how they perceive the impact of these initiatives in shaping the future of technology in their field (see Table 11). The participants could respond on a Likert Scale with an option to describe the impact in more detail. The majority (66.5%) sees at least a somewhat positive impact of the initiatives on the future of technology. However, many participants were uncertain or neutral (23.1%). About 10% consider a negative impact. One participant noted: “We already have that in Europe. I’m all hears for data and privacy protection laws, but sometimes these laws are maded by people outside of the industry and this has had a negative impact on the European Tech Market before.” About half of those who responded yes are mostly concerned that regulations may limit innovation. One participant remarked: “I believe that regulations for mitigating AI ethics risks might initially pose challenges, such as increased compliance costs and slower innovation. However, they can ultimately benefit the industry by building trust, ensuring fair practices, and fostering a more responsible and sustainable approach to AI development.” Others are concerned about operational challenges and costs associated with regulations. According to one participant: “If the company fails to comply with the regulations, it may face legal action, fines or other legal consequences. This not only causes direct financial losses, but may also damage the company’s reputation.” Among those who do not believe regulation negatively impacts their work, a common theme is the belief that regulation can enhance trust across the AI ecosystem. In the words of one participant: “As said, we need to ensure the data protection before using AI for our daily business. Having a higher authority to regulate that safety will increase our and our client’s trust in the systems.” Among those who are unsure, we observe a balance between recognizing the positive aspects of regulations and concerns about limiting innovation and competitiveness. One respondent commented: ”Regulations for mitigating AI ethics risks might increase compliance costs and slow innovation, potentially creating competitive disadvantages for smaller companies. However its also important to maintain certain standards.”
We conducted a LASSO analysis to identify predictive factors (using the same approach as in the previous analysis) most strongly associated with participants’ perceived impact of AI regulations, ranging from highly negative to highly positive. For Group B, our model shows that more positive perceptions of regulatory impact are most strongly associated with how frequently participants consider AI ethics principles during development (+0.333) and with greater familiarity with environmental and social responsibility (+0.284) (see Table 12).
Table 12: LASSO Results: Predictors of Regulatory Impact Perception (Gr. B)
Predictor Coefficient Std. Error
AI Practice
   Frequency of considering ethics in development $0.333^{**}$ $(0.109)$
Familiarity with AI Ethics Principles
   Environment and Social Responsibility $0.284^{**}$ $(0.105)$
Overall Mean 3.68
Observations 137
R² (test) 0.10
Features Selected 2 / 150
Note: $^{*}p<0.05$; $^{**}p<0.01$; $^{***}p<0.001$. Dependent Var.: Perceived impact of AI regulations (1–5 scale, highly -ive to highly +ive). Coefficients are standardized (per 1-SD change in predictor).
@@
RQ2 & RQ3 - AI Governance Initiatives: Overall, the average familiarity with the AI governance initiatives is lower than the familiarity with the principles among various roles and demographics. However, most AI development teams have a positive outlook on how regulations will impact the future of AI development and their companies. The field is expected to benefit from increased user trust fostered by regulations and their associated safety guidelines. However, many express concern about the effects of these regulations, particularly regarding the challenges of implementing guidelines into AI technology, which may result in slowing innovation and reducing competitiveness.
4.3.5 Predictors of AI Ethics Principle Familiarity and Practice
Here, we present the results of linear mixed-effects models used to identify statistically significant predictors of familiarity and practice related to AI ethics principles and initiatives (i.e., RQ2 and RQ3).
Table 13: Fixed Effects in Linear Mixed-Effects Models
Variable Effect SE $z$ $p$-value
Model 1: Principle Familiarity ($n=267$, $k=6$ roles)
Intercept 2.465 0.164 15.03 <0.001<0.001
Experience 0.074 0.045 1.64 0.101
Develops AI 0.056-0.056 0.114 0.49-0.49 0.623
Governance Initiative Familiarity 0.419 0.036 11.65 <0.001<0.001
Model 2: Ethics Consideration Frequency ($n=267$, $k=6$ roles)
Intercept 2.009 0.309 6.50 <0.001<0.001
Experience 0.019 0.057 0.34 0.735
Principle Familiarity 0.296 0.078 3.80 <0.001<0.001
Governance Initiative Familiarity 0.238 0.056 4.23 <0.001<0.001
Predictors of AI Ethics Principle Familiarity
We evaluate which factors predict familiarity with AI ethics principles (i.e., Model 1). Model 1 reveals that familiarity with AI governance initiatives is the only statistically significant predictor ($\beta=0.419$, $p<0.001$), whereas years of experience in AI development is not a significant predictor ($\beta=0.074$, $p=0.101$). Each 1-point increase in governance initiative familiarity predicts a 0.419-point increase in principle familiarity, indicating that knowledge of initiatives and principles is strongly linked.
Predictors for Considering AI Ethics Principles
We also examine which factors predict the frequency of considering AI ethics principles (i.e., Model 2). Model 2 shows that both principle familiarity ($\beta=0.296$, $SE=0.078$, $z=3.80$, $p<0.001$) and governance initiative familiarity ($\beta=0.238$, $SE=0.056$, $z=4.23$, $p<0.001$) are significant predictors, while experience is non-significant ($\beta=0.019$, $SE=0.057$, $p=0.735$). The likelihood ratio tests confirmed that both types of familiarity independently improve model fit: adding governance initiative familiarity to a model of only principles ($\chi^{2}(1)=12.45$, $p<0.001$) and adding principle familiarity to a model of only familiarity with AI governance initiatives ($\chi^{2}(1)=10.69$, $p=0.001$) both significantly improved fit. The model that includes both familiarity types ($AIC=701.34$) fit substantially better than models with only principles ($AIC=711.80$) or only initiatives ($AIC=710.03$).
Hypotheses Testing
We conduct hypothesis tests to analyze the impact of various factors on familiarity and consideration of AI ethics principles.
H4a: Does AI development experience predict familiarity with AI ethics principles? The experience effect is $\beta=0.182$, $SE=0.054$, $z=3.37$, corresponding to $p=0.001$. We reject H4a_0H4a\_0 (no effect) at $\alpha=0.05$. Each additional experience level predicts a 0.182-point increase in familiarity with principles. A practitioner with 10+ years scores 0.55 points higher than one with 1–2 years—a small but statistically significant effect.
H4b: Does familiarity with AI governance initiatives predict frequency of considering AI ethics principles in work? The AI governance initiative familiarity effect is $\beta=0.348$, $SE=0.047$, $z=7.37$, corresponding to $p<0.001$. We strongly reject H4b_0H4b\_0 (no effect) at $\alpha=0.05$. Each 1-point increase in initiative familiarity predicts a 0.348-point increase in ethics consideration frequency—a moderate to large effect.
H4c: Does professional role predict familiarity and consideration of AI ethics? The Intraclass Correlation Coefficient (ICC) quantifies variance due to role clustering. We got $ICC=0.029$ (2.9% between roles) and $ICC=0.155$ (15.5% between roles) for AI ethics principle familiarity and AI ethics frequency, respectively. We partially reject H4c_0H4c\_0, suggesting that role has a negligible impact on knowledge but a moderate impact on practice frequency.
Model Selection and Interaction Effects
To validate our findings regarding the significance of fixed effects, we compared nested models using Likelihood Ratio Tests (LRT) with significance level $\alpha=0.05$.
[…]
Table 14: Model Comparison via Likelihood Ratio Tests
[…]
Principle Familiarity Models ($n=271$)
Null Intercept only 717.60
M1a Experience 705.48 12.12-12.12 12.04 0.001
M1b Experience + Develops AI 706.27 +0.79+0.79 1.20 0.271
M1c Experience $\times$ Location + Develops AI 698.03 8.24-8.24 2.45-2.45 1.000
M1d Experience $\times$ Develops AI + Location 689.99 8.04-8.04 3.60 0.309
Ethics Frequency Models ($n=272$)
Null Intercept only 788.82
M2a Governance Initiative familiarity 736.36 52.46-52.46 52.03 <0.001<0.001
M2b Reg. fam. + Experience 737.13 +0.77+0.77 1.24 0.265
M2c Reg. fam. $\times$ Location + Experience 734.38 2.75-2.75 5.62-5.62 1.000
M2d Reg. fam. $\times$ Experience + Location 727.78 6.60-6.60 1.02-1.02 1.000
[…]
Effect Sizes
We calculated marginal $R^{2}$ and conditional $R^{2}$ to quantify the variance in the outcome that each model explains [nakagawa2013general]. Marginal $R^{2}$ represents variance explained by fixed effects alone (i.e., the predictors of interest), while conditional $R^{2}$ includes variance explained by both fixed and random effects (i.e., accounting for clustering within roles). The difference between them indicates the extent to which additional variance is attributable to role-based grouping. For Model 1, fixed effects explained 6.3% of variance (marginal $R^{2}=0.063$), increasing to 9.4% when accounting for role clustering (conditional $R^{2}=0.094$). This shows that practitioners within the same role tend to exhibit similar levels of familiarity with the principles. For Model 2, familiarity with initiatives explained 17.4% of variance (marginal $R^{2}=0.174$), increasing to 27.8% with role clustering (conditional $R^{2}=0.278$)—a moderate effect. The larger gap between the marginal and conditional $R^{2}$ suggests that professional roles play a substantial role in explaining variation in how frequently they consider AI ethics principles.
Interpretation
These findings reveal asymmetry: experience predicts knowledge ($\beta=0.182$, $p=0.001$) but accounts for only a small portion of variance (6.3%), whereas familiarity with governance initiative strongly predicts ethical practice ($\beta=0.348$, $p<0.001$) with moderate explained variance (17.4%). This indicates that simply having more years of experience does not necessarily lead to more frequent consideration of ethics; rather, specific knowledge of AI governance initiatives drives ethical practice. Role-based clustering is minimal for knowledge (ICC = 2.9%) but moderate for practice (ICC = 15.5%), suggesting that organizational and role-specific norms influence behavior more than individual knowledge acquisition. These effects remain consistent across locations and experience levels.
RQ2 & RQ3 - Predictive Factors: Linear mixed-effects modeling revealed that AI development experience significantly predicts familiarity with ethics principles, whereas familiarity with AI governance initiatives is the strongest predictor of how frequently practitioners consider ethics in their work. A professional role has a minimal impact on knowledge but has a meaningful influence on ethical practice. Interaction effects were tested but found to be non-significant, indicating that these relationships are consistent across locations and experience levels. These findings suggest that targeted initiatives, education, and role-specific interventions may be more effective than experience-based training alone in promoting ethical AI development practices.
4.4 Risk Mitigation Strategies
@@
In RQ4, our goal is to understand the actual practices employed by members of AI development teams to mitigate AI ethics risks. We tailored questions on AI ethics risk mitigation strategies to each role within Group B, as detailed in Table 1.
@@
4.4.1 AI Manager and Administrative (AM) Roles
Those in AM roles in Group B rated how effectively AI ethics principles are integrated into their products, with 57.38% reporting this as “Extremely” or “Moderately Effective,” and only 13.12% rating it as “Moderately Ineffective” or worse.
Table 15: Effectiveness of AI Ethics Principles Integration
Extremely Effective Moderately Effective Somewhat Effective Moderately Ineffective Somewhat Ineffective Prefer not to say
27.87% 29.51% 27.87% 8.20% 4.92% 1.64%
@@ 4.4.1 AI Manager and Administrative (AM) Roles
Figure 9: Group B - Risk Mitigation Methods Used
Table 16: P-Value and Chi-Square Statistic for Hypotheses H5a to H5e
H5a H5b H5c H5d H5e
@@
We followed up by asking those in AM roles to identify the actions their company is taking to promote AI ethics. Table 17 shows the results. “Providing AI ethics training for employees” and “implementing ethics guidelines for AI development and deployment” are considered key actions. However, only $\sim$50% of the AM roles conduct regular audits of the AI systems for ethics compliance, which is concerning given that one of the mitigation strategies in OWASP Top 10 for LLMs and AI models [owasp2025llm] is to “implement robust logging and monitoring systems to detect unusual patterns.”
Breaking down these responses by location, we observe that participants from North America are most likely to establish an AI ethics committee (57.14%), while those from the EU+EEA+UK are more likely to implement ethics guidelines (63.64%). In contrast, participants outside of these two regions take other actions, e.g., “conduct regular audit AI systems for ethical compliance” (86.7%), “provide AI ethics training for employees” (73.3%), and/or “engage with external experts on AI ethics” (73.3%).
They then selected who is responsible for ensuring AI ethics principles are followed. The results are shown in Table 18 (note that participants were allowed to select multiple options): 63.93% pointed to the AI development team and upper management, while only 42.62% pointed to the internal legal and compliance team or the AI ethics committee. Previous research has highlighted the critical role of dedicated teams in ensuring software/AI compliance [prybylo2024evaluating]. Other studies have emphasized the challenges of enforcing accountability for AI ethics and regulatory standards and the importance of collaborating with policymakers to implement them [orr2020attributions].
Table 17: AM Roles: Actions Taken at Company to Promote AI Ethics
@@
Table 18: AM Roles: Primary Responsibility for Following AI Ethics Principles
@@
Lastly, we asked them to provide examples of how they ensure their team adheres to AI ethics principles in development and describe how they promote and enforce them. Four mentioned “developing ethical guidelines,” while one participant mentioned they have been “developing AI-specific red-teams.” Although the responses about promoting and enforcing the principles varied, we observed that 30.3%, 21.21%, and 21.21% of participants mentioned implementing ethical guidelines, communicating AI guidelines with employees, and educating employees on AI ethics, respectively. Of those who mentioned communicating with employees, participants highlighted, among others: “understanding and teaching staff that AI is there to aid the system, not BE the system” and “use of stringent measures on failure to adopt the principles.” We also gained insights into some of the challenges faced during this process. For instance, one participant noted: “our role mainly revolves around building what the client requires, and we cannot control the end usage.” Another added, “we are a small company still with less than 5 workers so to keep the information up to date is easy right now.”
Table 19: Frequency of Ethical Consideration in AI Requirements Documentation
Always Often Sometimes Rarely Never
Percentage 20.6% 35.3% 20.6% 17.6% 5.9%
4.4.2 Requirements Analyst (RA) Roles
Those in RA roles described how often they include AI ethics principles in their requirements documentation. Interestingly, fewer than 25% of them rarely or never consider ethics in their requirements documentation (see Table 19). Similarly, $\sim$30% believe that AI ethics requirements hinder, to some degree, the project’s outcome (see Table 20). Regarding whether their company conducts (ethics-related) risk and impact assessments of their products, $\sim$53% of responses were positive, with one participant noting that “another unit is responsible.” The rest responded “no” or “unsure.”